The difference of two numbers is 63. The second is 9 less than 4 times the first. What are the two numbers?
Answer by ankor@dixie-net.com(22740)

The difference of two numbers is 63. The second is 9 less than 4 times the first. What are the two numbers?
:
Let x = a number:
:
Just write an equation/expression for what it says:
"The difference of two numbers is 63"
x and (x + 63)
:
"The second is 9 less than 4 times the first."
(x + 63) = 4x - 9
:
What are the two numbers?
x + 63 = 4x - 9
:
63 + 9 = 4x - x
:
72 = 3x
:
x = 72/3
:
x = 24
:
Our two numbers are 24 and 87
:
Check our answers in the statement:
"The second is 9 less than 4 times the first."
87 = 4(24) - 9
87 = 96 - 9; confirms our solution